OpenStudy (sparkyz):

Finding surface area of an irregular shape given the volume 52,000,000^3

OpenStudy (sparkyz):

I just want to know how I would find the answer

OpenStudy (tkhunny):

Not enough information. You can give it an interesting guess. Volume of a Sphere: \(\dfrac{4}{3}\pi\;r^{3}\) Surface Area of a Shpere: \(4\pi\;r^{2}\) Use the first to create a very rough approximation of a regular "r" and use THAT to produce the very rough approximation of the surface area.